# Leveraging Priors via Diffusion Bridge for Time Series Generation

### v1

@@ Appendix C Experimental settings
Appendix D Additional Experimental Results
Generation with longer input
We conduct experiments to compare the long-term generation performance of our model with Diffusion-TS (Yuan and Qiao 2024), specifically over a time length of 64. The experimental results are shown in Table 11. We reproduced Diffusion-TS and achieved better performance compared to their results using a small number of selected data samples. Except for the Energy dataset, the proposed TimeBridge-GP outperformed in generation quality with an input length of 64.
Table 11: Quality comparison of unconditional time series synthetic data with a length of 64.
Metric Methods Datasets
ETTh Stocks Energy fMRI
Context-FID score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.235±.011 0.948±.227 0.067±.010 3.662±.104
TimeBridge-GP 0.110±.009 0.693±.165 0.072±.010 0.527±.024
Correlational score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.061±.010 0.005±.005 0.442±.077 4.585±.068
TimeBridge-GP 0.038±.011 0.005±.003 0.548±.102 1.569±.028
Discriminative score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.078±.017 0.149±.035 0.065±.011 0.245±.234
TimeBridge-GP 0.045±.013 0.090±.018 0.149±.005 0.211±.184
Predictive score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.115±.002 0.038±.001 0.249±.010 0.103±.000
TimeBridge-GP 0.114±.005 0.038±.000 0.250±.001 0.100±.000
Generation with same NFE
Table 12 presents the experimental results for Diffusion-TS (Yuan and Qiao 2024) with the same NFE of 119 across most datasets. The results demonstrate that TimeBridge outperforms Diffusion-TS by using better priors for time series and the second-order sampler, compared to standard Gaussian priors.
Table 12: Comparison of generation performance with same NFE of 119.
@@ Appendix D Additional Experimental Results
Context-FID score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.014±.001 0.033±.006 0.244±.012 0.252±.014 0.126±.026 0.149±.009
@@
Correlational score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.015±.002 0.194±.023 0.071±.011 0.021±.004 0.878±.144 1.079±.024
@@
Discriminative score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.043±.014 0.040±.008 0.103±.006 0.130±.009 0.143±.008 0.156±.016
@@
Predictive score (Lower is better) Diffusion-TS (Yuan and Qiao 2024) 0.095±.000 0.008±.001 0.125±.005 0.038±.000 0.251±.000 0.099±.000
@@
Predictive score
Predictive score (Yoon, Jarrett, and Van der Schaar 2019) measures the performance to predict the next step trained with synthetic data as a train-synthesis-and-test-real (TSTR) method. Table 13 illustrates the predictive score in the same experimental setting of Table 1. Similarly, Table 14 shows the predictive score of the same experiments as in Table 2.
Table 13: Predictive score unconditional time series synthetic data. The best results are in bold and the second best are underlined.
@@
Predictive score (Lower is better) TimeVAE (Desai et al. 2021) 0.093±.000 0.012±.002 0.126±.004 0.039±.000 0.292±.000 0.113±.003
TimeGAN (Yoon, Jarrett, and Van der Schaar 2019) 0.093±.019 0.025±.003 0.124±.001 0.038±.001 0.273±.004 0.126±.002
Cot-GAN (Xu et al. 2020) 0.100±.000 0.068±.009 0.129±.000 0.047±.001 0.259±.000 0.185±.003
Diffwave (Kong et al. 2021) 0.093±.000 0.013±.000 0.130±.001 0.047±.000 0.251±.000 0.101±.000
Diffusion-TS (Yuan and Qiao 2024) 0.093±.000 0.007±.001 0.123±.003 0.037±.000 0.251±.000 0.100±.000
@@
The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao 2024). We re-implement Diffusion-TS with their official code.
Table 14: Quality comparison of trend-conditioned time series synthetic data.
@@
SSSD (Alcaraz and Strodthoff 2023) 0.1228±.0041 0.1206±.0036 0.1214±.0047 0.2498±.0005 0.2501±.0004 0.2498±.0004
Diffusion-TS (Yuan and Qiao 2024) 0.1204±.0055 0.1200±.0059 0.1197±.0060 0.2503±.0003 0.2500±.0005 0.2495±.0007
@@
Ablation study
VP and VE scheduler.
@@
Table 15: Quality comparison of unconditional time series with VP and VE scheduler.
@@
Importance of $D$-matching in time series.
@@
Table 16: Quality comparison of unconditional time series with $D$-matching and $F$-matching.
@@
Table 17: Quality comparison of unconditional time series with sampling step ratio $s$.
@@
Fourier-based loss.
@@
Table 18: Quality comparison of unconditional time series without Fourier.
@@
Appendix E Visualization
Visualization of distribution properties.
In addition to visualization of the synthetic data on t-SNE Figure 3, we visualize the synthetic data of TimeBridge using PCA kernel and kernel density estimation in Figures 7 and 8, respectively. More visualization results on baseline methods can be found in the Appendix of Diffusion-TS (Yuan and Qiao 2024).
@@ Appendix E Visualization
Figure 7: Visualization on PCA. Red and blue indicate original and synthetic data, respectively.
@@
Figure 8: Visualization of the kernel density estimation. Red and blue indicate original and synthetic data, respectively.
Sampling paths.
@@
(a) Unconditional (GP)
(b) Linear trend
(c) Polynomial trend
(d) Butterworth trend
Figure 9: Sampling path from (Left) the prior to (Right) the data samples with diverse prior settings on the ETTh dataset.
Figure 10: Sampling path from the prior (Left) to the imputation (Right) of TimeBridge (red) using point-preserving sampling based on observed values (dots) of the truth (green).
@@
Figure 11: Visualization of imputation on (Top) Energy and (Bottom) ETTh dataset. The green dotted lines denote the true and the green circles denote the observed value. Diffusion-TS and TimeBridge are visualized with grey and red along with their interval.

### v2

@@ B.4. Additional Experimental Details
Details for sampling from GP Prior
We sample ${\mathbf{x}}_{T}$ in Equation (20) as follows:
(34) ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{m}(\Omega),\bm{\Sigma}(\Omega)),\quad\Omega=\{1,\ldots,\tau\}.$
Specifically, we set $\bm{m}(\Omega)$ and $\bm{\Sigma}(\Omega)$ by $\bm{m}(\Omega)=\bm{\mu}$ and $\bm{\Sigma}(\Omega)_{i,j}=\texttt{diag}(\bm{\sigma}^{2})_{i,j}+\exp\left(-\frac{|i-j|^{2}}{2l^{2}}\right)$ for timestamps $i$ and $j$ in Equation (19). $l$ is the normalization parameter using data standard deviation to encode data information in kernel function.
Appendix C Additional Experimental Results
C.1. Wasserstein distance
To experimentally justify the strength of the diffusion prior, we measure the Wasserstein distance within the diffusion bridge’s optimal transport framework. Kollovieh et al. (2024) argue that closer alignment of data and prior distributions reduces path complexity, and time-reflective priors can lower the Wasserstein distance. As shown in Table 13, our data-dependent priors (unconditional and imputation) yield shorter distances. While not always a direct indicator of performance, this may explain why diffusion bridge models surpass standard diffusion approaches.
Table 13. Wasserstein distance results for unconditional and imputation settings using different priors.
Unconditional Imputation
Prior $\mathcal{N}(\mathbf{0},\mathbf{I})$ $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ (Ours) $\mathcal{N}(\mathbf{0},\mathbf{I})$ Linear Spline (Ours)
ETTh 14.285 2.387 8.000 0.485
Energy 28.128 6.376 16.113 1.494
C.2. Extreme cases
To evaluate effectiveness on extreme values, we perform an ablation study on tailed data samples using real-world data in our experiments, i.e., ETTh, Stocks, and Energy datasets. For each real and synthetic dataset, we extract top $k\%$ data samples with high variance as extreme data, respectively. The results from samples with the top $20\%$ and $50\%$ variance are shown in Table 14. Our method outperformed 10 out of 12 metrics, showing its robustness in high variance and tailed data.
Table 14. Ablation results on extreme values of top $k\%$ variance.
Top $k\%$ Metric ETTh Energy Stock
Diffusion-TS Ours Diffusion-TS Ours Diffusion-TS Ours
20% C.-FID 0.153 0.103 0.128 0.093 0.788 0.257
Corr. 0.095 0.070 0.946 0.150 1.217 0.010
50% C.-FID 0.144 0.098 0.139 0.096 0.475 0.154
Corr. 0.062 0.048 0.886 0.021 1.028 0.012
C.3. Longer Sequences
Modeling longer sequences is crucial for time series generation. Thus, we have tested lengths of 192 and 360, common for forecasting cyclic or annual patterns. By controlling only $sigma_{d}ata$ and fixing other parameters to minimize the search space, Table 15 shows that our method outperforms standard diffusion priors, excelling in 11 out of 12 metrics. These results will be included in the main paper for real-world usage.
Table 15. Comparison of longer sequence lengths (192, 360). Lower is better.
Context-FID Cross-Correlation Discriminative
Dataset (length) Diffusion-TS Ours Diffusion-TS Ours Diffusion-TS Ours
Stocks (192) 0.679 0.548 0.009 0.006 0.144 0.060
Stocks (360) 0.954 0.606 0.016 0.004 0.125 0.069
ETTh (192) 1.021 1.535 0.106 0.098 0.187 0.096
ETTh (360) 1.977 1.780 0.114 0.099 0.226 0.109
C.4. Generation with same NFE
Table 16 presents the experimental results for Diffusion-TS (Yuan and Qiao, 2024) with the same NFE of 119 across most datasets. The results demonstrate that TimeBridge outperforms Diffusion-TS by using better priors for time series and the second-order sampler, compared to standard Gaussian priors.
Table 16. Comparison of generation performance with same NFE of 119.
@@ C.4. Generation with same NFE
Context-FID score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.014±.001 0.033±.006 0.244±.012 0.252±.014 0.126±.026 0.149±.009
@@
Correlational score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.015±.002 0.194±.023 0.071±.011 0.021±.004 0.878±.144 1.079±.024
@@
Discriminative score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.043±.014 0.040±.008 0.103±.006 0.130±.009 0.143±.008 0.156±.016
@@
Predictive score (Lower is better) Diffusion-TS (Yuan and Qiao, 2024) 0.095±.000 0.008±.001 0.125±.005 0.038±.000 0.251±.000 0.099±.000
@@
C.5. Predictive score
Predictive score (Yoon et al., 2019) measures the performance to predict the next step trained with synthetic data as a train-synthesis-and-test-real (TSTR) method. Table 17 illustrates the predictive score in the same experimental setting of Table 2. Similarly, Table 18 shows the predictive score of the same experiments as in Table 4.
Table 17. Predictive score unconditional time series synthetic data. The best results are in bold and the second best are underlined. The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao, 2024). We re-implement Diffusion-TS with their official code.
@@ C.5. Predictive score
Predictive score (Lower is better) TimeVAE (Desai et al., 2021) 0.093±.000 0.012±.002 0.126±.004 0.039±.000 0.292±.000 0.113±.003
TimeGAN (Yoon et al., 2019) 0.093±.019 0.025±.003 0.124±.001 0.038±.001 0.273±.004 0.126±.002
Cot-GAN (Xu et al., 2020) 0.100±.000 0.068±.009 0.129±.000 0.047±.001 0.259±.000 0.185±.003
Diffwave (Kong et al., 2021) 0.093±.000 0.013±.000 0.130±.001 0.047±.000 0.251±.000 0.101±.000
Diffusion-TS (Yuan and Qiao, 2024) 0.093±.000 0.007±.001 0.123±.003 0.037±.000 0.251±.000 0.100±.000
@@
Table 18. Quality comparison of trend-conditioned time series synthetic data.
@@
SSSD (Alcaraz and Strodthoff, 2023) 0.1228±.0041 0.1206±.0036 0.1214±.0047 0.2498±.0005 0.2501±.0004 0.2498±.0004
Diffusion-TS (Yuan and Qiao, 2024) 0.1204±.0055 0.1200±.0059 0.1197±.0060 0.2503±.0003 0.2500±.0005 0.2495±.0007
@@
C.6. Ablation study
VP and VE scheduler
@@ C.6. Ablation study
Table 19. Quality comparison of unconditional time series with VP and VE scheduler.
@@
Importance of $D$-matching in time series
@@
Table 20. Quality comparison of unconditional time series with $D$-matching and $F$-matching.
@@
Table 21. Quality comparison of unconditional time series with sampling step ratio $s$.
@@
Fourier-based loss
@@
Table 22. Quality comparison of unconditional time series without Fourier.
@@
C.7. Visualization
@@
Figure 7. Visualization on PCA. Red and blue indicate original and synthetic data, respectively.
Visualization of distribution properties
In addition to visualization of the synthetic data on t-SNE Figure 3, we visualize the synthetic data of TimeBridge using PCA kernel and kernel density estimation in Figures 7 and 8, respectively. More visualization results on baseline methods can be found in the Appendix of Diffusion-TS (Yuan and Qiao, 2024).
@@
Figure 8. Visualization of the kernel density estimation. Red and blue indicate original and synthetic data, respectively.
Sampling paths
@@
(a) Unconditional (GP) - First
(b) Unconditional (GP) - Second
(c) Linear Trend - First
(d) Linear Trend - Second
(e) Polynomial Trend - First
(f) Polynomial Trend - Second
(g) Butterworth Trend - First
(h) Butterworth Trend - Second
Figure 9. Sampling path from (Left) the prior to (Right) the data samples with diverse prior settings on the ETTh dataset.
Figure 10. Sampling path from the prior (Left) to the imputation (Right) of TimeBridge (red) using point-preserving sampling based on observed values (dots) of the truth (green).
@@
Figure 11. Visualization of imputation on (Top) Energy and (Bottom) ETTh datasets. The green dotted lines denote the true values, and the green circles denote the observed values. Diffusion-TS and TimeBridge are visualized with grey and red, along with their interval.